# DAG-Math: Graph-Guided Mathematical Reasoning in LLMs

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate strong performance on mathematical problems when prompted with Chain-of-Thought (CoT), yet it remains unclear whether this success stems from search, rote procedures, or rule-consistent reasoning. To address this, we propose modeling CoT as a certain rule-based stochastic process over directed acyclic graphs (DAGs), where nodes represent intermediate derivation states and edges encode rule applications. Within this framework, we introduce logical closeness, a metric that quantifies how well a model’s CoT trajectory (i.e., the LLM’s final output) adheres to the DAG structure, providing evaluation beyond classical PASS@$k$ metrics. Building on this, we introduce the DAG-MATH CoT format and construct a benchmark that guides LLMs to generate CoT trajectories in this format, thereby enabling the evaluation of their reasoning ability under our framework. Across standard mathematical reasoning datasets, our analysis uncovers statistically significant differences in reasoning fidelity among representative LLM families—even when PASS@$k$ is comparable—highlighting gaps between final-answer accuracy and rule-consistent derivation. Our framework provides a balance between free-form CoT and formal proofs systems, offering actionable diagnostics for LLMs reasoning evaluation. Our benchmark and code are available at https://github.com/YuanheZ/DAG-MATH-Formatted-CoT.
@@ 1 Introduction
The black-box nature of CoT in LLMs raises a key challenge: how to rigorously model and evaluate LLMs’ mathematical reasoning abilities. Prior works formalize reasoning in terms of search (Shalev-Shwartz & Shashua, 2025), probabilistic inference (Prystawski et al., 2023; Feng et al., 2023), and propositional logic (Merrill & Sabharwal, 2023; Kim & Suzuki, 2024; Yin et al., 2025). Intuitively, LLM reasoning needs to identify all required premises (e.g., facts, constraints), and conduct correct logical inference from premises to reach the conclusion. [^1] Accurate calculation and symbolic execution are also required, see the discussion in Section A.1. These operations must be exact, in line with the exact learning requirements in György et al. (2025). To test whether LLMs achieve this through CoT, two elements are crucial:
@@
Despite recent progress, existing approaches remain limited in both framework and evaluation. In terms of frameworks, prior work has examined Boolean-circuit analyses (Li et al., 2024), $k$-parity models (Kim & Suzuki, 2024; Yin et al., 2025), and graph-based representations (Dziri et al., 2023). Graph-based strategies provide a natural way to model CoT as a discrete, step-level, graph-based abstractions where nodes correspond to intermediate assertions and edges encode inferential dependencies. However, existing graph-based formulations, whether modeling CoT as deterministic subgraph matching in a directed acyclic graph (DAG) (Dziri et al., 2023), random walks on reversible/recurrent Markov chains (Kim et al., 2025), or trees of linear solution paths (Shalev-Shwartz & Shashua, 2025), neglect the improvement from diverse sampling (Wang et al., 2023) and ability to connect disparate knowledge (Yin & Wang, 2025). As a result, they fail to capture long-range and cross-branch dependencies, as well as the goal-directed, absorbing-state nature of CoT.
In terms of evaluations, prior work (Dziri et al., 2023; Joshi et al., 2025; Kim et al., 2025; Xu & Sato, 2025) primarily relies on final-answer metrics such as PASS@$k$ while the entire output for inference is overlooked, which leaves it unclear whether it is solved by logical inference or by search. A promising alternative is to use the LEAN programming language (De Moura et al., 2015; Moura & Ullrich, 2021) to formally verify solutions (Google DeepMind, 2024; Ren et al., 2025; Wang et al., 2025; Lin et al., 2025). While LEAN-based verification ensures logical correctness, it presupposes that problems have been already formalized in a proof-oriented form. For answer-based problems, the formalization into Lean must be carried out in advance, which requires substantial expert effort.
@@ 2 A DAG Framework for Step-Level CoT
Motivated by empirical observations in Bogdan et al. (2025), we study CoT at the step level, rather than the token level. This step-level perspective has been widely considered in recent theoretical analyses (Dziri et al., 2023; Hu et al., 2024; Kim et al., 2025; Shalev-Shwartz & Shashua, 2025), as it better captures intermediate reasoning and the logical structure of solutions. We model step-level CoT in a two-phase workflow as below. Phase 1 defines a task-specific DAG, where Phase 2 samples CoT trajectories over this DAG under certain stochastic process. For better illustration, we take the following problem, adapted from MATH-500 (Hendrycks et al., 2021), as a representative example.
@@ 2.1 Phase 1: Task-specific DAG for Step-Level CoT
Note that such a decomposition is non-unique due to semantic variation, such as synonyms or equivalent phrasings. This ambiguity makes it challenging to develop a precise and principled definition of a CoT step. Consequently, prior work (Dziri et al., 2023; Hu et al., 2024; Kim et al., 2025; Shalev-Shwartz & Shashua, 2025; Bogdan et al., 2025) has typically defined steps heuristically—either as text spans or task-specific annotations—without providing a formal definition. As a first attempt, we present an abstract mathematical formulation, with the technical details deferred to Section˜A.2 since they are not essential for understanding the main text. Within this formulation, although the node/edge decomposition for a single step may still be non-unique, the task-specific DAG introduced later can be made unique, provided that each step is restricted to a single conclusion.
@@
This assumption covers the answer-based math problems which have tractable computation graphs such as AIME (Art of Problem Solving, 2025a, b). Note that, if the correct sink node is included in ${\mathcal{G}}({\bm{x}}_{\tt in})$, the task-specific DAG can be always constructed by backtracking through its ancestors. We remark that, some “thinking-LLMs”, e.g., DeepSeek-R1 (DeepSeek Team, 2025), encourage backtracking and self-correction during the thinking process. However, we expect that LLMs output only the finalized, perfect, correct reasoning results to the user. The reasoning evaluation in this paper is based on the entire final output (while PASS@$k$ just considers the final answer). The definition of perfect and correct reasoning will be introduced in the next section.
@@ 2.2 Phase 2: Stochastic Process on Logic Dependence
Comparison with previous work: Our framework integrates an instance-specific DAG with a rule-based stochastic process, directly addressing key limitations in prior work. We do not assume a fixed “universal" graph with deterministic matching Dziri et al. (2023). Instead, our DAG is logical, acyclic, and features absorbing sink nodes, preserving the directional, goal-oriented nature of mathematical problem solving while allowing for long-range dependencies, in contrast to reversible Markov chain models (Kim et al., 2025). Furthermore, unlike the tree abstractions in Shalev-Shwartz & Shashua (2025), our DAG captures shared sub-derivations and the true dependency structure rather than just linear solution paths, thereby supporting multiple valid routes to a solution.
@@ 3 Formal Definition of Mathematical Reasoning Ability
Under Assumption˜2.1, consider an input prompt ${\bm{x}}_{\tt in}$ and the DAG ${\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})$. For each node $v\in{\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})$, define its out-degree as
@@
Under Assumption˜2.1, let an LLM be given a prompt ${\bm{X}}_{\tt in}\in{\mathcal{P}}$, sampled from an underlying distribution $\mathcal{D}$} over mathematical problem prompts. We define two indicator functions for a trajectory-specific DAG ${\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})$:
@@
To remain logically closed, a trajectory must stay on the same chain across the remaining L1L-1 transitions, each occurring with probability $1/2$. Hence, we have $\operatorname{PRR}=\left(\frac{1}{2}\right)^{L-1}.$ This illustrates that PRR decays exponentially with depth: although the final-answer accuracy ($1/2$) may appear stable, logically closed trajectories become increasingly rare. Furthermore, Appendix˜B presents an example from AIME 2025 (Art of Problem Solving, 2025a, b) where the final answer is correct, but logic-closeness fails due to mixing two solution paths. Consequently, lacking logic closeness risks producing answers correct only by chance, obscuring flawed reasoning, reducing reliability, and undermining interpretability.
@@ 4.2 Gold-Standard DAG-MATH Benchmark
The benchmark comprises problems from Omni-MATH (Gao et al., 2024), which are categorized into difficulty levels ranging from 1 (easiest) to 10 (hardest). To ensure high solvability by LLMs, we only consider problems with difficulty levels below 6. For generating DAG-MATH formatted CoTs, we employ GPT-o4-mini and Qwen3-235B-A22B-Thinking-2507, both recognized as leading models in mathematical tasks (LMArena, 2025). Gold-standard CoTs are constructed using a three-stage prompting strategy (see Appendix˜D) in reverse order (Node $\rightarrow$ Parents $\rightarrow$ Edge). This approach fixes the node set first, making verification easier with SymPy or LLM-as-Judge and minimizing error propagation, thereby ensuring high-quality trajectories.
@@ 5 Evaluation of Mathematical Reasoning Ability
To evaluate mathematical reasoning performance, we employ few-shot prompting (using demonstrations from the benchmark in Section˜4.2; see Appendix˜E for details) to guide LLMs in generating DAG-MATH formatted CoTs/DAGs for test problems, without providing the final solutions.
@@
Fig.˜4 reports AUC scores across three datasets, with PASS@1 as the starting point and $\widehat{\mathcal{R}}$ as the end point (see more details in Table˜2 in Appendix˜F), as the logic closeness rate increases. Besides, we also report averaged graph-level statistics for AIME 2025 in Table˜1, with additional results for BRUMO 2025 and HMMT 2025 in Appendix˜F. We have the following observations:
Figure 4: The AUC curves of averaged accuracy under the logical closeness rate over three datasets for five selected LLMs.
@@ A.2 Mathematical Definition of Steps in CoT
Appendix B Example of Logical Closeness
@@
[scale=0.045, line join=round, line cap=round, >=stealth]

### v2

# DAG-Math: Graph-of-Thought Guided Mathematical Reasoning in LLMs

@@ Abstract
Large Language Models (LLMs) demonstrate strong performance on mathematical problems when prompted with Chain-of-Thought (CoT), yet it remains unclear whether this success stems from search, rote procedures, or rule-consistent reasoning. To address this, we propose modeling CoT as a certain rule-based stochastic process over directed acyclic graphs (DAGs), where nodes represent intermediate derivation states and edges encode rule applications. Within this framework, we introduce logical closeness, a metric that quantifies how well a model’s CoT trajectory (i.e., the LLM’s final output) adheres to the DAG structure, providing evaluation beyond classical PASS@$k$ metrics. Building on this, we introduce the DAG-MATH CoT format and construct a benchmark that guides LLMs to generate CoT trajectories in this format, thereby enabling the evaluation of their reasoning ability under our framework. Across standard mathematical reasoning datasets, our analysis uncovers statistically significant differences in reasoning fidelity among representative LLM families—even when PASS@$k$ is comparable—highlighting gaps between final-answer accuracy and rule-consistent derivation. Our framework provides a balance between free-form CoT and formal proofs systems, offering actionable diagnostics for LLMs reasoning evaluation. Our benchmark and code are available at https://github.com/YuanheZ/DAG-MATH.
@@ 1 Introduction
The black-box nature of CoT in LLMs raises a key challenge: how to rigorously model and evaluate LLMs’ mathematical reasoning abilities. Prior works formalize reasoning in terms of search (Shalev-Shwartz and Shashua, 2025), probabilistic inference (Prystawski et al., 2023; Feng et al., 2023), and propositional logic (Merrill and Sabharwal, 2023; Kim and Suzuki, 2024; Yin et al., 2025). Intuitively, LLM reasoning needs to identify all required premises (e.g., facts, constraints), and conduct correct logical inference from premises to reach the conclusion. [^1] Accurate calculation and symbolic execution are also required, see the discussion in Section A.1. These operations must be exact, in line with the exact learning requirements in György et al. (2025). To test whether LLMs achieve this through CoT, two elements are crucial:
@@
Despite recent progress, existing approaches remain limited in both framework and evaluation. In terms of frameworks, prior work has examined Boolean-circuit analyses (Li et al., 2024), $k$-parity models (Kim and Suzuki, 2024; Yin et al., 2025), and graph-based representations (Dziri et al., 2023). Graph-based strategies provide a natural way to model CoT as a discrete, step-level, graph-based abstractions where nodes correspond to intermediate assertions and edges encode inferential dependencies. However, existing graph-based formulations, whether modeling CoT as deterministic subgraph matching in a directed acyclic graph (DAG) (Dziri et al., 2023), random walks on reversible/recurrent Markov chains (Kim et al., 2025), or trees of linear solution paths (Shalev-Shwartz and Shashua, 2025), neglect the improvement from diverse sampling (Wang et al., 2023) and ability to connect disparate knowledge (Yin and Wang, 2025). As a result, they fail to capture long-range and cross-branch dependencies, as well as the goal-directed, absorbing-state nature of CoT.
In terms of evaluations, prior work (Dziri et al., 2023; Joshi et al., 2025; Kim et al., 2025; Xu and Sato, 2025) primarily relies on final-answer metrics such as PASS@$k$ while the entire output for inference is overlooked, which leaves it unclear whether it is solved by logical inference or by search. A promising alternative is to use the LEAN programming language (De Moura et al., 2015; Moura and Ullrich, 2021) to formally verify solutions (Google DeepMind, 2024; Ren et al., 2025; Wang et al., 2025; Lin et al., 2025). While LEAN-based verification ensures logical correctness, it presupposes that problems have been already formalized in a proof-oriented form. For answer-based problems, the formalization into Lean must be carried out in advance, which requires substantial expert effort.
@@ 2 A DAG Framework for Step-Level CoT
Motivated by empirical observations in Bogdan et al. (2025), we study CoT at the step level, rather than the token level. This step-level perspective has been widely considered in recent theoretical analyses (Dziri et al., 2023; Hu et al., 2024; Kim et al., 2025; Shalev-Shwartz and Shashua, 2025), as it better captures intermediate reasoning and the logical structure of solutions. We model step-level CoT in a two-phase workflow as below. Phase 1 defines a task-specific DAG, where Phase 2 samples CoT trajectories over this DAG under certain stochastic process. For better illustration, we take the following problem, adapted from MATH-500 (Hendrycks et al., 2021), as a representative example.
@@ 2.1 Phase 1: Task-specific DAG for Step-Level CoT
Note that such a decomposition is non-unique due to semantic variation, such as synonyms or equivalent phrasings. This ambiguity makes it challenging to develop a precise and principled definition of a CoT step. Consequently, prior work (Dziri et al., 2023; Hu et al., 2024; Kim et al., 2025; Shalev-Shwartz and Shashua, 2025; Bogdan et al., 2025) has typically defined steps heuristically—either as text spans or task-specific annotations—without providing a formal definition. As a first attempt, we present an abstract mathematical formulation, with the technical details deferred to Section˜A.2 since they are not essential for understanding the main text. Within this formulation, although the node/edge decomposition for a single step may still be non-unique, the task-specific DAG introduced later can be made unique, provided that each step is restricted to a single conclusion.
@@
If the correct sink node is included in ${\mathcal{G}}({\bm{x}}_{\tt in})$, the task-specific DAG can be always constructed by backtracking through its ancestors. Note that, the LLM may reason imperfectly during its thinking process, e.g., dead-ends, self-correction, but is expected to output only the finalized, perfect, correct reasoning results to the user (with formal definition later). The reasoning evaluation in this paper is based on the entire final output while PASS@$k$ just considers the final answer. More discussion can be found in Section˜A.3.
@@ 2.2 Phase 2: Stochastic Process on Logic Dependence
Comparison with previous work: Our framework integrates an instance-specific DAG with a rule-based stochastic process, directly addressing key limitations in prior work. We do not assume a fixed “universal" graph with deterministic matching Dziri et al. (2023). Instead, our DAG is logical, acyclic, and features absorbing sink nodes, preserving the directional, goal-oriented nature of mathematical problem solving while allowing for long-range dependencies, in contrast to reversible Markov chain models (Kim et al., 2025). Furthermore, unlike the tree abstractions in Shalev-Shwartz and Shashua (2025), our DAG captures shared sub-derivations and the true dependency structure rather than just linear solution paths, thereby supporting multiple valid routes to a solution.
@@ 3 Formal Definition of Mathematical Reasoning Ability
Under Assumption˜1, consider an input prompt ${\bm{x}}_{\tt in}$ and the DAG ${\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})$. For each node $v\in{\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})$, define its out-degree as
@@
Under Assumption˜1, let an LLM be given a prompt ${\bm{X}}_{\tt in}\in{\mathcal{P}}$, sampled from an underlying distribution $\mathcal{D}$} over mathematical problem prompts. We define two indicator functions for a trajectory-specific DAG ${\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})$:
@@
To remain logically closed, a trajectory must stay on the same chain across the remaining L1L-1 transitions, each occurring with probability $1/2$. Hence, we have $\operatorname{PRR}=\left(\frac{1}{2}\right)^{L-1}.$ This illustrates that PRR decays exponentially with depth: although the final-answer accuracy ($1/2$) may appear stable, logically closed trajectories become increasingly rare. Furthermore, Section˜B.1 presents a geometry example from AIME 2025 (Art of Problem Solving, 2025a) where the final answer is correct, but logic-closeness fails due to mixing two solution paths. Consequently, lacking logic closeness risks producing answers correct only by chance, obscuring flawed reasoning, reducing reliability, and undermining interpretability. Furthermore, our framework also can be extended to analyze IMO-level problems via lemma-level DAG, which can reveal the key structural distinction. A concrete example based on the problem 4 from IMO 2025 (IMO Foundation, 2025) is presented in Section˜B.2.
@@ 4.2 Gold-Standard DAG-MATH Benchmark
The benchmark comprises problems from Omni-MATH (Gao et al., 2024), which are categorized into difficulty levels ranging from 1 (easiest) to 10 (hardest). To ensure high solvability by LLMs, we only consider problems with difficulty levels below 6. For generating DAG-MATH formatted CoTs, we employ GPT-o4-mini and Qwen3-235B-A22B-Thinking-2507, both recognized as leading models in mathematical tasks (LMArena, 2025). Gold-standard CoTs are constructed using a three-stage prompting strategy (see Appendix˜D) in reverse order (Node $\rightarrow$ Parents $\rightarrow$ Edge). This approach fixes the node set first, making verification easier with SymPy or LLM-as-Judge and minimizing error propagation, thereby ensuring high-quality trajectories. Furthermore, we perform human evaluation on 50 random samples and the results (49/50) confirm the reliability of our benchmark, see more details in Appendix˜E.
@@ 5 Evaluation of Mathematical Reasoning Ability
To evaluate mathematical reasoning performance, we employ few-shot prompting (using demonstrations from the benchmark in Section˜4.2; see Appendix˜F for details) to guide LLMs in generating DAG-MATH formatted CoTs/DAGs for test problems, without providing the final solutions.
@@
Fig.˜4 reports AUC scores across three datasets, with PASS@1 as the starting point and $\widehat{\mathcal{R}}$ as the end point (see more details in Table˜4 in Appendix˜I), as the logic closeness rate increases. Besides, we also report averaged graph-level statistics for AIME 2025 in Table˜1, with additional results for BRUMO 2025 and HMMT 2025 in Appendix˜I. We have the following observations:
Figure 4: The AUC curves of averaged accuracy under the logical closeness rate over three datasets for five selected LLMs. The curve is a direct plot of the computed accuracies at a fine-grained grids of threshold (0.01 increment from 0 to 1).
@@
These insightful findings from our proposed metrics can provide a principled foundation for future work aimed at improvement, see a detailed discussion in Appendix˜J.
5.1 More experiments and discussions
Here we discuss several additional experiments to validate our claims and findings.
Reasoning Gap: We expand evaluation sample size from 32 to 128 trajectories per problem and calculate 95% confidence intervals for the gap between PASS@1 and PRR. The resulting intervals were tight and consistently excluded zero (e.g., confirming a  36% gap for Gemini-2.5-Flash), statistically demonstrating that the difference between finding the correct answer and generating a perfectly valid derivation is significant and substantial across all tested models. More details are presented in Section˜K.1.
Prompt Sensitivity: We employ Two One-Sided Tests (TOST) to verify that DAG-MATH evaluations are robust to variations in prompt. By perturbing the few-shot prompts through re-formatting (changing nested bullets to tables) and re-phrasing (using synonyms and alternative grammar), then generate new sets of DAGs. The analysis shows statistical equivalence across five key graph-level statistics and performance metrics, see more details in Section˜K.2.
Parsing Consistency and Non-Circularity: We validate the objectivity of PRR/AUC metric by examining the consistency of DAG construction across different parsing methods, including self-parsing and external parsing by Thinking LLMs. The AUC scores for logical closeness remained remarkably stable across all models (ranging from 23.8% to 25.5%) with statistically equivalent DAG structures tested by TOST, demonstrating that logical closeness is an intrinsic property of the solution rather than an artifact of the parser. More details are reported in Section˜K.3.
Formatting Constraints: By employing external models to parse natural CoT into dependency graphs, the analysis revealed that the metrics effectively capture fundamental reasoning patterns, with Natural CoT achieving AUC scores comparable to the structured format and over 70% logical closeness. This confirms that the evaluation framework is not limited to the benchmark but serves as a robust, objective method for diagnosing the logical coherence and reasoning fidelity of freeform LLM generations. More details are presented in Section˜K.3.
Cross-Family Few-Shot Demonstrations: We investigate potential biases introduced by the model source of few-shot demonstrations. While using cross-family examples resulted in a slight performance decrease (GPT-4.1 PRR drops from 16.8% to 15.9%), the DAG structures are statistically equivalent and the relative rankings of the models remained unchanged. This indicates that while same-family demonstrations are optimal, our framework’s evaluations remain valid even when cross-family prompting is employed. More details are presented in Section˜K.4.
Table 2: Results of PASS@1 and PRR for Gemini models w/wo thinking on AIME 2025.
PASS@1 PRR
Thinking w/o w w/o w
Flash 52.4% 64.0% 17.0% 35.9%
Flash-Lite 37.4% 49.2% 15.9% 24.5%
Role of Thinking: We perform experiments with thinking modes for two Gemini models om AIME 2025. The results in Table˜2 confirm that while thinking mode significantly boosts both PASS@1 and PRR, the substantial performance gap between two metrics persist. The continued gap indicates that thinking enhances the model’s exploration of the task-specific DAG, yielding more correct and logically closed trajectories, but does not eliminate the tendency on search over logical coherence.
@@ A.2 Mathematical Definition of Steps in CoT
A.3 Discussions on Acyclicity
[…]
Appendix B Examples of Logical Closeness
B.1 A High-School Competition-Level Example DAG
@@
[scale=0.035, line join=round, line cap=round, >=stealth]
@@
B.2 An IMO-Level Example DAG
There are several ways in which an IMO-level proof can fail the logical-closeness test even when substantial portions of the argument are correct:
  • a model may solve the fixed-point subproblem and mistake fixed points for all valid starting values;
  • a model may prove the forward dynamics for $N\in\mathcal{S}$ but omit the complementary necessity argument for $N\notin\mathcal{S}$;
  • a model may derive both macro-branches but never fuse them into a single final characterization.
For such problems, a lemma-level DAG is more informative than a step-level DAG. We illustrate this with the following problem from IMO 2025.
IMO 2025 Q4 Problem A proper divisor of a positive integer NN is a positive divisor of NN other than NN itself. The infinite sequence $a_{1},a_{2},\ldots$ consists of positive integers, each of which has at least three proper divisors. For each $n\geq 1$, the integer $a_{n+1}$ is the sum of the three largest proper divisors of $a_{n}$. Determine all possible values of $a_{1}$.
The rigorous solution shows that the valid starting values are exactly
$\mathcal{S}=\{N:\ v_{2}(N)\text{ is odd},\ v_{3}(N)\geq\tfrac{v_{2}(N)+1}{2},\ 5\nmid N\}.$
The resulting lemma-level task-specific DAG drawn from Claude-Opus-4-6 is shown in Fig.˜7.
Figure 7: Lemma-level task-specific DAG for IMO 2025 Q4. The green sink is the correct theorem, while the red sink is a plausible but incorrect premature termination.
We summarize each node’s conclusion below.
1
Defines the dynamical map $f(N)$, validity of a starting value, and the candidate set $\mathcal{S}$.
2
Proves the divisor-pairing identity $f(N)=N\bigl(\frac{1}{s_{2}}+\frac{1}{s_{3}}+\frac{1}{s_{4}}\bigr)$.
3
Solves $\frac{1}{a}+\frac{1}{b}+\frac{1}{c}=1$ in strictly increasing integers and obtains the unique triple $(2,3,6)$.
4
Characterizes the fixed points: $f(N)=N$ if and only if $N=2\cdot 3^{b}\cdot m$ with $b\geq 1$ and $\gcd(m,30)=1$.
5
Shows that odd inputs stay odd and strictly decrease under $f$.
6
Shows that when $12\mid N$ and $5\nmid N$, one has $f(N)=\frac{13N}{12}$ and $(v_{2},v_{3})\mapsto(v_{2}-2,v_{3}-1)$.
7
Proves sufficiency: every $N\in\mathcal{S}$ is valid.
8
Proves forward invariance of the complement: $N\notin\mathcal{S}\Rightarrow f(N)\notin\mathcal{S}$.
9
Proves necessity: every $N\notin\mathcal{S}$ eventually becomes invalid.
L
Correct sink: the valid starting values are exactly the elements of $\mathcal{S}$.
L
Incorrect sink: the valid starting values are exactly the fixed points of $f$.
The green sink requires both macro-branches of the proof. On the left, Nodes 2–4–6–7 show that numbers in $\mathcal{S}$ are not necessarily fixed at time 0, but repeated use of Node 6 decreases $v_{2}$ by $2$ and $v_{3}$ by $1$ until the orbit reaches the fixed-point family from Node 4. Concretely, if
$N=2^{2k+1}\cdot 3^{b}\cdot m\in\mathcal{S}\qquad(b\geq k+1,\ \gcd(m,30)=1),$
then after $k$ applications one obtains
$f^{k}(N)=\left(\frac{13}{12}\right)^{k}N=2\cdot 3^{b-k}\cdot 13^{k}m,$
which lies in the fixed-point family and therefore generates an infinite valid orbit.
On the right, Nodes 5, 6, 8, and 9 show that the complement of $\mathcal{S}$ cannot support an infinite orbit. Node 8 keeps every iterate outside $\mathcal{S}$. Then Node 9 selects the minimum element of the orbit and proves that any local growth pattern must fall into either the short-lived $(2,3,5)$ regime or the $(2,3,4)$ regime; both eventually force strict descent, contradicting well-foundedness of $\mathbb{N}$.
The incorrect sink $L^{\prime}$ models a natural but wrong termination. After Nodes 2, 3, and 4, a model might conclude that the only infinite sequences are the fixed points of $f$. This is a coherent subproof, but it misses the genuinely valid preperiodic starting values in $\mathcal{S}$ with $v_{2}(N)\geq 3$, which are not fixed at time 0 and only flow to a fixed point after finitely many applications of Node 6.
This failure mode is especially useful for evaluating logical closeness. A trajectory ending at $L^{\prime}$ has solved a substantial subproblem-the complete classification of fixed points-yet it omits the lemmas that distinguish fixed points from all valid starting values. In other words, $L^{\prime}$ is not a random wrong answer; it is a mathematically polished but globally incomplete termination.
At the DAG level, the proof therefore separates into a left sufficiency branch and a right necessity branch, both of which are indispensable for the correct sink. A logically closed trajectory must collect enough ancestors to justify Node L, rather than stopping as soon as Node 4 is established. For IMO-level arguments of this type, the lemma-level DAG exposes the key structural distinction much more clearly than a long list of atomic algebraic steps.